OpenStudy (jaina1537):

Solve the following system of equations and show all work. y = 2x^2 y = –3x −1

OpenStudy (aussiebabe23):

The two equations equal the same thing (y) so you can set them equal to each other to solve for x

OpenStudy (jaina1537):

How? I tried to but I got stuck.

OpenStudy (anthonyym):

This is the transitive property/ substitution. If A = B A = C, then B = C. You can apply this to your problem.

OpenStudy (aussiebabe23):

You should get 2x^2 + 3x + 1 = 0

OpenStudy (jaina1537):

I got to that part. I just don't know how to go further.

OpenStudy (anthonyym):

Now it just comes down to factoring. You need two numbers that add to the coefficient of the x term and multiply to the product of the constant and coefficient of x^2. So two numbers that add to 3 and multiply to 2.

OpenStudy (jaina1537):

So 2 and 1? (2x+1)(x+1)

OpenStudy (anthonyym):

Yes that's the correct factorization. (2x+1)(x+1) = 0 Now notice you have (2x+1) multiplied by (x+1) equals 0, so either 2x+1 or x+1 = 0 by the zero product property. So, 2x+1 = 0 or x + 1 = 0